# AI-accelerated Discovery of Altermagnetic Materials

### v3

@@ Abstract
Altermagnetism, a new magnetic phase, has been theoretically proposed and experimentally verified to be distinct from ferromagnetism and antiferromagnetism. Although altermagnets have been found to possess many exotic physical properties, the limited availability of known altermagnetic materials hinders the study of such properties. Hence, discovering more types of altermagnetic materials with different properties is crucial for a comprehensive understanding of altermagnetism and thus facilitating new applications in the next generation information technologies, e.g., storage devices and high-sensitivity sensors. Since each altermagnetic material has a unique crystal structure, we propose an automated discovery approach empowered by an AI search engine that employs a pre-trained graph neural network to learn the intrinsic features of the material crystal structure, followed by fine-tuning a classifier with limited positive samples to predict the altermagnetism probability of a given material candidate. Finally, we successfully discovered 50 new altermagnetic materials that cover metals, semiconductors, and insulators confirmed by the first-principles electronic structure calculations. The wide range of electronic structural characteristics reveals that various novel physical properties manifest in these newly discovered altermagnetic materials, e.g., anomalous Hall effect, anomalous Kerr effect, and topological property. Noteworthy, we discovered 4 $i$-wave altermagnetic materials for the first time. Overall, the AI search engine performs much better than human experts and suggests a set of new altermagnetic materials with unique properties, outlining its potential for accelerated discovery of the materials with targeted properties.
Keywords: altermagnetism, pre-trained model, symmetry analysis, material discovery
@@ Introduction
Magnetic materials form a cornerstone of our modern information society. Generally, magnetism is categorized into ferromagnetism and antiferromagnetism. Recently, based on the spin group formalism, a new magnetic phase called altermagnetism has been theoretically proposed [1, 2], which exhibits numerous novel physical properties [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], paving the path way of new avenues in the next generation of information technology. Both altermagnets and conventional antiferromagnets have compensated antiparallel spin sublattices resulting in vanishing net magnetic momentum. The compensated antiparallel spin sublattices are connected by the spin symmetry $\{C_{2}^{\perp}||I\}$ or $\{C_{2}^{\perp}||\tau\}$ transformation for conventional antiferromagnets, but by the spin symmetry $\{C_{2}^{\perp}||R_{i}\}$ transformation for altermagnets [1]. Here, the symmetry operations at the left and right of the double vertical bar act only on the spin space and lattice space, respectively; the notation $C_{2}^{\perp}$ represents the $180^{\circ}$ rotation perpendicular to the spin direction; the notations $I,T,R_{i}$, and $\tau$ denote space inversion, time reversal, rotation/mirror, and fractional translation operations, respectively. Due to the absence of spin symmetry $\{C_{2}^{\perp}||I\}$ or $\{C_{2}^{\perp}||\tau\}$, altermagnets have spin splitting in electronic bands. Unlike isotropic k-independent $s$-wave spin splitting in ferromagnets, altermagnets can form anisotropic k-dependent $d$-wave, $g$-wave, and $i$-wave spin splitting according to different spin group symmetry [1]. Moreover, altermagnets have not only spin-splitting bands deriving from magnetic exchange interaction which is the same as ferromagnets but also unique extraordinary spin-splitting bands deriving from anisotropic electric crystal potential and magnetic exchange interaction [1]. In some altermagnets, the spin splitting can even have magnitudes of $eV$ in parts of the Brillouin zone (BZ) [1, 2, 3]. The anisotropic k-dependent spin splitting can result in a unique spin current by electrical means in $d$-wave altermagnets [4]. Based on the unique spin current, the spin-splitter torque in $d$-wave altermagnets was proposed in theory [4] and confirmed by experiments [5, 6], which may circumvent limitations of spin-transfer torque (ferromagnets) and spin-orbit torque (conventional antiferromagnets or nonmagnetic materials with strong spin-orbit coupling) in magnetic memory devices [4]. Meanwhile, the giant, tunneling magnetoresistance [3] and giant piezomagnetisme [15] can also be proposed in altermagnets based on the anisotropic k-dependent spin splitting. In the relativistic case, the time-reversal symmetry-breaking macroscopic phenomena, including quantum anomalous Hall [7], anomalous Hall [8, 9], and anomalous Kerr effects [10], have been predicted by theories in altermagnets, moreover, the anomalous Hall effect has been supported by experiments [11, 12].
@@
On the other hand, magnetic topological phases and their exotic physical properties have recently attracted intensive experimental and theoretical attention. Very recently, some topological semimetal and insulator phases protected by spin group symmetry have been proposed in theory [16, 17, 18, 19]. Considering that the altermagnets are described by spin group symmetry and the symmetry landscape of spin space groups is more plentiful than that of the conventional magnetic space groups, more new magnetic topological phases and their exotic physical properties may be thus proposed theoretically in altermagnets. Nevertheless, altermagnets are hitherto in the early stage of research. Since there are many exotic physical properties that have been discovered and new physical phenomena to be discovered, altermagnets are bound to attract intensive theoretical and experimental attention in the near future. Very recently, based on spin group theory and known magnetic structures, 141 altermagnetic materials have been discovered [19]. However, altermagnetic materials are known still limited so far. Hence, there is an urgent need to discover more altermagnetic materials for a comprehensive understanding of altermagnetism, thus facilitating new applications in the next generation of information technologies.
@@
Conventional discovery methods primarily rely on the known magnetic structures and the corresponding spin space group. However, there exist over 90,000 magnetic materials documented in the Material Project [20], among which only 2,138 magnetic structures are known (see the MAGNDATA database [21]). The reason why the magnetic structures of only about 2% of magnetic materials have been determined is that it is indeed a non-trivial task that relies on extremely costly neutron scattering experimentation. On the other hand, the altermagnetic property is closely related to the material crystal structure, which provides a basis for the application of artificial intelligence (AI) methods to the discovery of altermagnetic materials. Moreover, the emerging AI technology has found many key applications in the discovery of materials [22]. For instance, AI was used for predicting organic compound synthesis in organic chemistry [23], planning chemical synthesis pathways [24], iterative synthesis of small molecules [25], accelerating the discovery of self-assembling peptides [26], designing eutectic solvent [27], and analyzing de novo protein mechanics and sturctures [28, 29]. Recently, deep learning methods have been applied to the prediction of crystal materials with targeted properties [30, 31]. These methods generally utilize a large amount of crystal structure data to train graph neural network (GNN) models in an end-to-end manner, without explicit reference to the physical laws underlying these material properties. The trained model could predict key physical properties of crystal materials, such as formation energy and band gap, based on a rich training dataset containing over $10^{4}$ labeled samples [30]. However, such methods are no longer suitable for discovering altermagnetic materials, because of the fact that the known positive samples are limited.
@@
In this article, we introduce an AI search engine, as shown in Fig. 1, that combines deep model pre-training and fine-tuning techniques and physics-based approaches (e.g., symmetry analysis and first-principles electronic structure calculation) to discover new altermagnetic materials under the condition of limited labeled samples. In particular, we pre-train a self-supervised GNN [32] based on optimal transport theory [33] to learn the intrinsic features of the crystal structure of materials, and refine a downstream classifier with limited positive samples to predict the altermagnetism probability of a given material candidate. First, based on symmetry analysis, we constructed the pre-training dataset (containing 68,116 materials), fine-tuning dataset (containing 25,739 materials, namely, 25,591 negative samples plus 148 positive samples), and candidate dataset (containing 42,377 materials) from the Material Project [34]. Next, we pre-trained a GNN model (composed of an encoder and a decoder) for crystal materials, based on optimal transport theory. Once the pre-training was done, we fine-tuned the encoder on the fine-tuning dataset to obtain a classifier model. Then, the structured material information from the candidate dataset was input into the classifier model to quantify the probability as an indicator of whether each material is an altermagnetic material. We filtered out the materials with probabilities greater than 0.9 as the candidate altermagnetic materials. Finally, we employed the first-principles electronic structure calculations to obtain the ground magnetic structure of the candidate material to identified altermagnets. Furthermore, the confirmed altermagnetic materials were added to the fine-tuning dataset for an iterative process of fine-tuning and classifier prediction, reinforcing the predictability of the model. The efficacy of this AI search engine has been well demonstrated.
@@
Of 91,649 total candidates, we discovered 50 new altermagnetic materials covering metals, semiconductors, and insulators. The wide range of electronic structural properties implies that various novel physical properties appear in these newly discovered altermagnetic materials, e.g., anomalous Hall effect, anomalous Kerr effect, and topological property, as demonstrated in theoretical analyses. It is also worth noting that we discovered 4 $i$-wave altermagnetic materials for the first time, filling in the gap in the literature. We conclude that the AI search engine performs much better than human experts and suggests a set of new altermagnetic materials with unique properties, outlining its potential for accelerated discovery of the materials with targeting properties. We also discuss the pathway of developing pre-trained graph models for the discovery of other types of materials.
@@ Results
Dataset screening based on symmetry analysis
Our goal is to screen altermagnetic materials from the Material Project [20], which contains 154,718 crystal materials. Since this material database includes both magnetic and non-magnetic materials, we first filtered out materials containing magnetic atoms. In this work, we considered materials with 3d transition metals or 4f rare earth elements. After filtering and de-duplication, we obtained 91,649 potential magnetic materials. Due to the complexity of the magnetic properties of materials with multiple magnetic atoms, we further excluded such materials, resulting in 68,116 potential magnetic materials, which constitute the pre-training dataset.
@@
Furthermore, magnetic materials with type-III magnetic space group symmetry can be divided into two classes according to space-inversion symmetry. If the crystal structure of a collinear antiferromagnet has no combination of space-inversion and time-reversal symmetry, such a material must be altermagnetic. Otherwise, if the collinear antiferromagnetic material has only a pair of spin antiparallel magnetic atoms in the primitive crystal cell that are not located at invariant space-inversion points, the pair of spin antiparallel magnetic atoms must be connected by the spin symmetry $\{C_{2}^{\perp}||I\}$. That is to say, the class collinear antiferromagnets are non-altermagnetic materials (7,045 in total). Therefore, based on symmetry analysis, we screened out 25,591 non-altermagnetic materials. These materials, along with the known 148 altermagnetic materials (e.g., as positive samples), together constitute the fine-tuning dataset. By removing the 25,591 non-altermagnetic materials and positive samples from the pre-training dataset, we obtained the candidate dataset (42,377 materials). The aforementioned screening process is depicted in Supplementary Fig. S.1. In the following, we train a neural network to screen and predict altermagnetic materials from the candidate dataset.
@@
As detailed in the symmetry analysis above, we first construct the pre-training dataset and candidate materials dataset based on high-throughput screening and symmetry analysis (see Fig. 1a). The pre-trained model is based on a GNN that leverages material crystal structure information  [30], consisting of a graph convolutional network encoder, and a decoder that reconstructs graph features based on the optimal transport theory [39]. Fig. 1b depicts the schematic of the network, with the detailed architecture shown in Fig. 2. The process of inputting crystal structures into the model begins with a pre-processing stage, where the crystal structure information is transformed into a graph representation. Then, we pre-train the model based on the pre-training dataset which contains 68,116 materials, and then fine-tune the pre-trained model based on the fine-tuning dataset (148 altermagnetic materials plus 25,591 non-altermagnetic materials) (see Fig. 1c). During the fine-tuning, we utilize the pre-trained encoder and employ up-sampling techniques (duplication and rotation) to balance the number of positive and negative samples for a binary classification task. Afterward, we can obtain the classifier model, which is then used to screen the altermagnetic materials (Fig. 1d). All possible candidate crystal structures (42,377) are input into the classifier model for prediction. The model provides a probability estimate for each sample, and we selected the material with a probability greater than 0.9 as the candidate material. Next, we utilize the first principle electronic structure calculations (Fig. 1e) to verify whether the candidates are altermagnetic materials. Furthermore, once the new altermagnetic materials are verified and confirmed, we added the new one to the fine-tuning dataset and then re-perform the fine-tuning and prediction iteratively. Through four rounds of iteration and leveraging information from 148 known altermagnetic materials, we identified 50 new altermagnetic materials. The discussion of the model convergence and the spin patterns distinguish ability is discussed in Supplementary Note C. The additional information for the pre-trained model is found in Supplementary Note A.
@@
To demonstrate the capability of our pre-trained crystal material model, we fed all the candidate materials in batches into the pre-trained encoder which provides a corresponding latent space vector for each material. We utilized principal component analysis (PCA) for dimensionality reduction and performed feature visualization and t-SNE visualization on the latent space vectors (see  Supplementary Fig. S.2). The results show that the data in the candidate set have a clear clustering phenomenon after pre-training, which indicates that the pre-training process can group materials containing similar information together.

### v4

# RESEARCH ARTICLE

@@ Abstract
[ABSTRACT]Altermagnetism, a new magnetic phase, has been theoretically proposed and experimentally verified to be distinct from ferromagnetism and antiferromagnetism. Although altermagnets have been found to possess many exotic physical properties, the limited availability of known altermagnetic materials hinders the study of such properties. Hence, discovering more types of altermagnetic materials with different properties is crucial for a comprehensive understanding of altermagnetism and thus facilitating new applications in the next generation of information technologies, e.g., storage devices and high-sensitivity sensors. Since each altermagnetic material has a unique crystal structure, we propose an automated discovery approach empowered by an AI search engine that employs a pre-trained graph neural network to learn the intrinsic features of the material crystal structure, followed by fine-tuning a classifier with limited positive samples to predict the altermagnetism probability of a given material candidate. Finally, we successfully discovered 50 new altermagnetic materials that cover metals, semiconductors, and insulators confirmed by the first-principles electronic structure calculations. The wide range of electronic structural characteristics reveals that various novel physical properties manifest in these newly discovered altermagnetic materials, e.g., anomalous Hall effect, anomalous Kerr effect, and topological property. Noteworthy, we discovered 4 $i$-wave altermagnetic materials for the first time. Overall, the AI search engine performs much better than human experts and suggests a set of new altermagnetic materials with unique properties, outlining its potential for accelerated discovery of the materials with targeted properties.
doi:
10.1093/nsr/nwaf066
keywords:
altermagnetism, pre-trained model, symmetry analysis, material discovery
@@ Introduction
Magnetic materials form a cornerstone of our modern information society. Generally, magnetism is categorized into ferromagnetism and antiferromagnetism. Recently, based on the spin group formalism [1], a new magnetic phase called altermagnetism has been theoretically proposed [2, 3], which exhibits numerous novel physical properties [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], paving the path way of new avenues in the next generation of information technology. Both altermagnets and conventional antiferromagnets have compensated antiparallel spin sublattices resulting in vanishing net magnetic momentum. The compensated antiparallel spin sublattices are connected by the spin symmetry $\{C_{2}^{\perp}||I\}$ or $\{C_{2}^{\perp}||\tau\}$ transformation for conventional antiferromagnets, but by the spin symmetry $\{C_{2}^{\perp}||R_{i}\}$ transformation for altermagnets [2]. Here, the symmetry operations at the left and right of the double vertical bar act only on the spin space and lattice space, respectively; the notation $C_{2}^{\perp}$ represents the $180^{\circ}$ rotation perpendicular to the spin direction; the notations $I,T,R_{i}$, and $\tau$ denote space inversion, time reversal, rotation/mirror, and fractional translation operations, respectively. Due to the absence of spin symmetry $\{C_{2}^{\perp}T||IT\}$ or $\{C_{2}^{\perp}||\tau\}$, altermagnets have spin splitting in electronic bands. Unlike isotropic k-independent $s$-wave spin splitting in ferromagnets, altermagnets can form anisotropic k-dependent $d$-wave, $g$-wave, and $i$-wave spin splitting according to different spin group symmetry [2]. Moreover, altermagnets have not only spin-splitting bands deriving from magnetic exchange interaction which is the same as ferromagnets but also unique extraordinary spin-splitting bands deriving from anisotropic electric crystal potential and magnetic exchange interaction [2]. In some altermagnets, the spin splitting can even have magnitudes of $eV$ in parts of the Brillouin zone (BZ) [2, 3, 4]. The anisotropic k-dependent spin splitting can result in a unique spin current by electrical means in $d$-wave altermagnets [5]. Based on the unique spin current, the spin-splitter torque in $d$-wave altermagnets was proposed in theory [5] and confirmed by experiments [6, 7], which may circumvent limitations of spin-transfer torque (ferromagnets) and spin-orbit torque (conventional antiferromagnets or nonmagnetic materials with strong spin-orbit coupling) in magnetic memory devices [5]. Meanwhile, the giant, tunneling magnetoresistance [4] and giant piezomagnetisme [19] can also be proposed in altermagnets based on the anisotropic k-dependent spin splitting. In the relativistic case, the time-reversal symmetry-breaking macroscopic phenomena, including quantum anomalous Hall [8], anomalous Hall [9, 10], and anomalous Kerr effects [11], have been predicted by theories in altermagnets, moreover, the anomalous Hall effect has been supported by experiments [12, 13].
@@
On the other hand, magnetic topological phases and their exotic physical properties have recently attracted intensive experimental and theoretical attention. Very recently, some topological semimetal and insulator phases protected by spin group symmetry have been proposed in theory [20, 21, 22, 23, 24, 25, 26]. Considering that the altermagnets are described by spin group symmetry and the symmetry landscape of spin space groups is more plentiful than that of the conventional magnetic space groups, more new magnetic topological phases and their exotic physical properties may be thus proposed theoretically in altermagnets. Nevertheless, altermagnets are hitherto in the early stage of research. Since there are many exotic physical properties that have been discovered and new physical phenomena to be discovered, altermagnets are bound to attract intensive theoretical and experimental attention in the near future. Very recently, based on spin group theory and known magnetic structures, 141 altermagnetic materials have been discovered [23]. However, known altermagnetic materials are still limited so far. Hence, there is an urgent need to discover more altermagnetic materials for a comprehensive understanding of altermagnetism, thus facilitating new applications in the next generation of information technology.
@@
Conventional discovery methods primarily rely on the known magnetic structures and the corresponding spin space group. Such approaches are applicable only when the magnetic structure information is known a priori, which has a clear limitation if such information is missing. However, there exist over 90,000 magnetic materials documented in the Material Project [jain2013materials], among which only 2,138 magnetic structures are known (see the MAGNDATA database [27]). The reason why the magnetic structures of only about 2% of magnetic materials have been determined is that it is indeed a non-trivial task that relies on extremely costly neutron scattering experimentation. Therefore, it is crucial to develop a method that breaks the bottleneck limitation of missing magnetic structure information, enabling the discovery of new altermagnetic materials without any prior knowledge of such information. On the other hand, the altermagnetic property is closely related to the material crystal structure, which provides a basis for the application of artificial intelligence (AI) methods to the discovery of altermagnetic materials. Moreover, the emerging AI technology has found many key applications in the discovery of materials [28]. For instance, AI was used for predicting organic compound synthesis in organic chemistry [29], planning chemical synthesis pathways [30], iterative synthesis of small molecules [31], accelerating the discovery of self-assembling peptides [32], designing eutectic solvent [33], and analyzing de novo protein mechanics and structures [34, 35]. Recently, deep learning methods have been applied to the prediction of crystal materials with targeted properties [36, 37]. These methods generally utilize a large amount of crystal structure data to train graph neural network (GNN) models in an end-to-end manner, without explicit reference to the physical laws underlying these material properties. The trained model could predict key physical properties of crystal materials, such as formation energy and band gap, based on a rich training dataset containing over $10^{4}$ labeled samples [36]. However, such methods are not suitable for discovering altermagnetic materials, because of the fact that the known positive samples are limited.
@@
In this article, we introduce an AI search engine, as shown in Fig. 1, that combines deep model pre-training and fine-tuning techniques and physics-based approaches (e.g., symmetry analysis and first-principles electronic structure calculation) to discover new altermagnetic materials under the condition of limited labeled samples. In particular, we pre-train a self-supervised GNN [38] based on optimal transport theory [39] to learn the intrinsic features of the crystal structure of materials, and refine a downstream classifier with limited positive samples to predict the altermagnetism probability of a given material candidate. First, based on symmetry analysis, we constructed the pre-training dataset (containing 68,116 materials), fine-tuning dataset (containing 25,739 materials, namely, 25,591 negative samples plus 148 positive samples), and candidate dataset (containing 42,377 materials) from the Material Project [40]. Next, we pre-trained a GNN model (composed of an encoder and a decoder) for crystal materials, based on optimal transport theory. Once the pre-training was done, we fine-tuned the encoder on the fine-tuning dataset to obtain a classifier model. Then, the structured material information from the candidate dataset was input into the classifier model to quantify the probability as an indicator of whether each material is an altermagnetic material. We filtered out the materials with probabilities greater than 0.9 as the candidate altermagnetic materials. Finally, we employed the first-principles electronic structure calculations to estimate the ground magnetic structure of the candidate material to identify altermagnets [^1] It has been quite common to use the first-principles electronic structure calculations, e.g., density functional theory (DFT), to predict the material property, which has been widely used in the community and proven to possess excellent alignment with experimental results for crystal materials [41, 42, 43, 44].. Furthermore, the confirmed altermagnetic materials were added to the fine-tuning dataset for an iterative process of fine-tuning and classifier prediction, reinforcing the predictability of the model. The efficacy of this AI search engine has been well demonstrated.
@@
Of 91,649 total candidates, we discovered 50 new altermagnetic materials covering metals, semiconductors, and insulators. The wide range of electronic structural properties implies that various novel physical properties appear in these newly discovered altermagnetic materials, e.g., anomalous Hall effect, anomalous Kerr effect, and topological property, as demonstrated in theoretical analyses. It is also worth noting that we discovered 4 $i$-wave altermagnetic materials for the first time, filling in the gap in the literature. As a result, our proposed AI search engine successfully breaks the bottleneck limitation of existing discovery methods based on symmetry delimited rules, serves as a critical counterpart to such methods, and is applicable to discovering new altermagnetic materials directly from a large set of candidates without any prior knowledge of the magnetic structure information. We conclude that the AI search engine suggests a set of new altermagnetic materials with unique properties, outlining its potential for accelerated discovery of the materials with targeting properties. We also discuss the pathway of developing pre-trained graph models for the discovery of other types of materials.
@@ Results
Figure 1: Workflow of the pre-trained model for searching altermagnetic materials. a, Construction of candidate material datasets using high-throughput screening and symmetry analysis (see Supplementary Appendix Fig. S.1 for details). b, The pre-training autoencoder framework for crystal materials. The input of the model is the crystal structure. Each crystal structure can be represented as a multi-edge graph neural network (GNN). The encoder is built by the graph convolutional neural network. The decoder is built on the Waterstein neighborhood reconstruction. c, The fine-tuning procedure with loading pre-training stage weight matrix. d, The prediction procedure by inputting candidate materials. e, Validation of the altermagnetic property via the first principle electronic structure calculations.
Dataset screening via symmetry analysis
Our goal is to screen altermagnetic materials from the Material Project [jain2013materials], which contains 154,718 crystal materials. Since this material database includes both magnetic and non-magnetic materials, we first filtered out materials containing magnetic atoms. In this work, we considered materials with 3d transition metals or 4f rare earth elements. After filtering and de-duplication, we obtained 91,649 potential magnetic materials. Due to the complexity of the magnetic properties of materials with multiple magnetic atoms, we further excluded such materials, resulting in 68,116 potential magnetic materials, which constitute the pre-training dataset.
@@
Furthermore, magnetic materials with type-III magnetic space group symmetry can be divided into two classes according to space-inversion symmetry. If the crystal structure of a collinear antiferromagnet has no combination of space-inversion and time-reversal symmetry, such a material must be altermagnetic. Otherwise, if the collinear antiferromagnetic material has only a pair of spin antiparallel magnetic atoms in the primitive crystal cell that are not located at invariant space-inversion points, the pair of spin antiparallel magnetic atoms must be connected by the spin symmetry $\{C_{2}^{\perp}||I\}$. That is to say, the class collinear antiferromagnets are non-altermagnetic materials (7,045 in total). Therefore, based on symmetry analysis, we screened out 25,591 non-altermagnetic materials. These materials, along with the known 148 altermagnetic materials (e.g., as positive samples), constitute the fine-tuning dataset. By removing the 25,591 non-altermagnetic materials and positive samples from the pre-training dataset, we obtained the candidate dataset (42,377 materials). The aforementioned screening process is depicted in Supplementary Appendix Fig. S.1. In the following, we train a neural network to screen and predict altermagnetic materials from the candidate dataset.
@@
As detailed in the symmetry analysis above, we first construct the pre-training dataset and candidate materials dataset based on high-throughput screening and symmetry analysis (see Fig. 1a). The pre-trained model is based on a GNN that leverages material crystal structure information  [36], consisting of a graph convolutional network encoder, and a decoder that reconstructs graph features based on the optimal transport theory [49]. Fig. 1b depicts the schematic of the network, with the detailed architecture shown in Fig. 2. The process of inputting crystal structures into the model begins with a pre-processing stage, where the crystal structure information is transformed into a graph representation. Then, we pre-train the model based on the pre-training dataset which contains 68,116 materials, and then fine-tune the pre-trained model based on the fine-tuning dataset (148 altermagnetic materials plus 25,591 non-altermagnetic materials  [^2] Note that there are some biases in the negative sampling, but its influence on the predictive performance of our AI model is negligible. This is because the number of negative samples is significantly larger, being 172 times greater than the number of positive samples.) (see Fig. 1c). During the fine-tuning, we utilize the pre-trained encoder and employ up-sampling techniques (duplication and rotation) to balance the number of positive and negative samples for a binary classification task. Afterward, we can obtain the classifier model, which is then used to screen the altermagnetic materials (Fig. 1d). All possible candidate crystal structures (42,377) are input into the classifier model for prediction. The model provides a probability estimate for each sample, and we selected the material with a probability greater than 0.9 as the candidate material. Next, we utilize the first principle electronic structure calculations (Fig. 1e) to verify whether the candidates are altermagnetic materials. Furthermore, once the new altermagnetic materials are verified and confirmed, we add the new one to the fine-tuning dataset and then re-perform the fine-tuning and prediction iteratively. Through four rounds of iteration and leveraging information from 148 known altermagnetic materials, we identified 50 new altermagnetic materials. The discussion of the model convergence and the spin patterns distinguish ability is discussed in Supplementary Appendix Note C. The additional information for the pre-trained model is found in Supplementary Appendix Note A.
@@
To demonstrate the capability of our pre-trained crystal material model, we fed all the candidate materials in batches into the pre-trained encoder which provides a corresponding latent space vector for each material. We utilized principal component analysis (PCA) for dimensionality reduction and performed feature visualization and t-SNE visualization on the latent space vectors (see Supplementary Appendix Fig. S.2). The results show that the data in the candidate set have a clear clustering phenomenon after pre-training, which indicates that the pre-training process can group materials containing similar information together.
1 $\text{Nb}_{2}\text{Fe}\text{B}_{2}$ $P4$/$mbm~{}(127)$ $g$-wave M mp-1086660 [51]
2 $\text{Ta}_{2}\text{Fe}\text{B}_{2}$ $P4$/$mbm~{}(127)$ $g$-wave M mp-1095076 [52]
3 $\text{Nd}\text{B}_{2}\text{C}_{2}$ $P4$/$mbm~{}(127)$ $g$-wave M mp-5765 [50]
4 $\text{Mg}_{2}\text{Fe}\text{Ir}_{5}\text{B}_{2}$ $P4$/$mbm~{}(127)$ $g$-wave M mp-1188243 [53]
5 $\text{Mg}_{2}\text{Mn}\text{Ir}_{5}\text{B}_{2}$ $P4$/$mbm~{}(127)$ $g$-wave M mp-1189623 [53]
6 $\text{Mg}_{2}\text{Ni}\text{Ir}_{5}\text{B}_{2}$ $P4$/$mbm~{}(127)$ $g$-wave M mp-1188248 [53]
7 $\text{Sc}_{2}\text{V}\text{Ir}_{5}\text{B}_{2}$ $P4$/$mbm~{}(127)$ $g$-wave M mp-20524 [54]
8 $\text{Sc}_{2}\text{Mn}\text{Ir}_{5}\text{B}_{2}$ $P4$/$mbm~{}(127)$ $g$-wave M mp-1208987 [54]
9 $\text{Ca}\text{La}\text{Fe}\text{Ag}\text{O}_{6}$ $Pc~{}(7)$ $d$-wave M mp-1641528 NA
10 $\text{Ca}\text{La}\text{Cr}_{2}\text{O}_{6}$ $Pmn2_{1}~{}(31)$ $d$-wave M mp-1642123 NA
11 $\text{Ni}\text{F}_{3}$ $R\bar{3}c~{}(167)$ $i$-wave M mp-561428 [55]
12 $\text{Gd}\text{B}_{2}\text{C}_{2}$ $P4/mbm~{}(127)$ $g$-wave M mp-1080176 [56]
13 $\text{Ho}\text{B}_{2}\text{C}_{2}$ $P4/mbm~{}(127)$ $g$-wave M mp-20410 [57]
14 $\text{Lu}\text{Cr}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave M mp-755471 [58]
15 $\text{Ta}\text{Co}\text{B}_{2}$ $Pnma~{}(62)$ $d$-wave M mp-1189690 NA
16 $\text{Nd}\text{Ru}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave M mp-1200843 [59]
17 $\text{Fe}\text{H}\text{O}_{2}$ $Pmn2_{1}~{}(31)$ $d$-wave I mp-510670 [60]
18 $\text{Na}\text{Fe}\text{O}_{2}$ $Pna2_{1}~{}(33)$ $d$-wave I mp-21060 [61]
19 $\text{Na}\text{Fe}\text{O}_{2}$ $P4_{1}2_{1}2~{}(92)$ $d$-wave I mp-21880 [62]
20 $\text{Mn}\text{O}_{2}$ $Pnma~{}(62)$ $d$-wave I mp-19326 [63]
21 $\text{Mn}\text{O}_{2}$ $I4/m~{}(87)$ $d$-wave I mp-19395 [64]
22 $\text{Ca}_{3}\text{Cr}_{2}\text{O}_{7}$ $Cmc2_{1}~{}(36)$ $d$-wave I mp-1575873 NA
23 $\text{Zr}\text{Cr}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave I mp-755055 NA
24 $\text{Zr}\text{Mn}\text{O}_{3}$ $R3c~{}(161)$ $i$-wave I mp-754513 NA
25 $\text{V}\text{F}_{3}$ $R\bar{3}c~{}(167)$ $i$-wave I mp-559931 [55]
26 $\text{Cr}\text{F}_{3}$ $R\bar{3}c~{}(167)$ $i$-wave I mp-560338 [55]
27 MnO $P6_{3}mc~{}(186)$ $g$-wave I mp-999539 [65]
28 $\text{Ca}\text{Mn}\text{N}_{2}$ $P6_{3}/mmc~{}(194)$ $g$-wave I mp-1246377 NA
29 $\text{Ba}_{2}\text{Fe}\text{Ge}_{2}\text{O}_{7}$ $P\overline{4}2_{1}m~{}(113)$ $g$-wave I mp-1190820 [66]
30 $\text{Ba}_{2}\text{Co}\text{Si}_{2}\text{O}_{7}$ $P\overline{4}2_{1}m~{}(113)$ $g$-wave I mp-510015 [67]
31 $\text{Sr}_{2}\text{Co}\text{Ge}_{2}\text{O}_{7}$ $P\overline{4}2_{1}m~{}(113)$ $g$-wave I mp-1191317 [68]
32 $\text{V}\text{F}_{4}$ $P2_{1}/c~{}(14)$ $d$-wave I mp-760030 NA
33 $\text{Ca}_{2}\text{Co}\text{Te}\text{O}_{6}$ $P2_{1}/c~{}(14)$ $d$-wave I mp-552051 [69]
34 $\text{Ni}\text{F}_{2}$ $Pnnm~{}(58)$ $d$-wave I mp-556324 [70]
35 $\text{Li}\text{Fe}_{2}\text{F}_{6}$ $P4_{2}nm~{}(102)$ $d$-wave I mp-557403 [71]
36 $\text{Fe}\text{H}\text{O}_{2}$ $P2_{1}2_{1}2_{1}~{}(19)$ $d$-wave I mp-625251 NA
37 $\text{Ca}\text{Mn}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave I mp-19201 [72]
38 $\text{Ca}\text{V}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave I mp-22608 [73]
39 $\text{La}\text{Fe}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave I mp-22590 [74]
40 $\text{La}\text{V}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave I mp-19350 [75]
41 $\text{Mn}\text{Se}\text{O}_{4}$ $Pnma~{}(62)$ $d$-wave I mp-817982 [76]
42 $\text{Na}\text{Pr}_{2}\text{Os}\text{O}_{6}$ $P2_{1}/c~{}(14)$ $d$-wave I mp-20009 [77]
43 $\text{Na}\text{Pr}_{2}\text{Ru}\text{O}_{6}$ $P2_{1}/c~{}(14)$ $d$-wave I mp-542512 [78]
44 $\text{Nd}\text{Rh}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave I mp-4582 [79]
45 $\text{Pr}\text{Ru}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave I mp-20186 [80]
46 $\text{Sc}\text{V}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave I mp-756546 [81]
47 $\text{Sm}\text{Rh}\text{O}_{3}$ $Pnma~{}(62)$ $d$-wave I mp-3317 [82]
48 $\text{Ca}\text{La}\text{Cr}\text{Mo}\text{O}_{6}$ $Pc~{}(7)$ $d$-wave I mp-1640189 NA
49 $\text{La}_{2}\text{Mn}\text{Rh}\text{O}_{6}$ $P2_{1}/c~{}(14)$ $d$-wave I mp-1223338 NA
50 $\text{Li}\text{Fe}\text{F}_{4}$ $P2_{1}/c~{}(14)$ $d$-wave I mp-755632 NA